A spring whose spring constant is 1000 N/m is compressed by 10 cm. By how much does its mass increase in its compressed state?
Answer #1

enegry due to compression= 1000*0.1*0.1/2

=5 j

using enegry mass equation, E=mc^2

M=5,56 *10^-17 kg